OpenStudy (anonymous):

a) are they linearly dependent? (-3,0,4),(5,-1,2),1,1,3) b)show that 3 vectors in R3 form a linearly dependent set in R4 v1=(o,3,1,-1); v2=(6,0,5,1);v3= (4,-7,1,3)

OpenStudy (anonymous):

I confused don't know when we use column/row to express the vectors in matrix to consider they are linearly dependent or not. Anyone explain me, please. just that stuff. the leftover is mine. I know how to prove it. Please.

OpenStudy (anonymous):

\[\left[\begin{matrix}-3 & 0 &4 \\ 5 & -1 & 2\\1 &1 & 3\end{matrix}\right]or \left[\begin{matrix}-3 & 5 &1 \\ 0 & -1 & 1\\4 & 2 & 3\end{matrix}\right]\]

OpenStudy (anonymous):

which one is the right matrix for me to consider they are linear dependent or not?

OpenStudy (anonymous):

@Hoa how can u unified 3 row vectors or row matrices in one matrix?

OpenStudy (anonymous):

it's the process to figure out whether they are linear dependent or independent. is it not right?

OpenStudy (anonymous):

no its not the process the process for determining linear dependence and independence of vectors is as follow : Vectors ( or matrices ) X1 , X2 , X3 ,... Xn are said to be dependent (1) all the vectors ( row or column ) are of the same order (2) n scalars a1 , a2 , a3, ..an(not all zero) exist such that a1X1 + a2X2 + a3X3+..+anXn = 0 otherwise they are linearly independent

OpenStudy (anonymous):

yes, the same idea, the same thing. I studied that information too. put your equation into the linear combination, you get exactly what i have. x1 = <-3,0,4> ...something like that. and you have the same matrix I have

OpenStudy (anonymous):

but u need to solve for the scalars in order to see whether they are linear independent or dependent

OpenStudy (anonymous):

you have 3 equations to solve for your a1,a2,a3 that means -3a1 +5a2+1a3=0 and then 0 a1-1a2+1a3 =0, and then 4a1 +2a2+3a3 =0 and you have exactly my matrix

OpenStudy (anonymous):

I think you confused. the matrix I set up is to solve for your scalar. if you have any scalar satisfy that matrix it means they are linear dependent. if not they are independent

OpenStudy (anonymous):

yeah i know u will get the second one matrix

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so row vectors are linearly dependent

OpenStudy (anonymous):

I don't solve it yet. just make sure about the matrix first, the leftover is easy

OpenStudy (anonymous):

u will get the second one matrix that for sure :-P

OpenStudy (anonymous):

i confused, because sometimes, my pro uses the first matrix, to me, the second one makes more sense than the first one.

OpenStudy (anonymous):

How can first matrix be used? I m confused too

OpenStudy (anonymous):

wait, i send email to ask my pro already. if i have his answer, i will let you know